OpenStudy (anonymous):

Order from greatest to least. 65 x 104, 432 x 103, 2.996 x 104 a. 432 x 103, 65 x 104, 2.996 x 104 b. 65 x 104, 432 x 103, 2.996 x 104 c. 2.996 x 104, 432 x 103, 65 x 104

OpenStudy (mathmate):

You can work this out quickly by approximations 65 x 104 ~ 65*100=6500 432 x 103 ~ 43*100=4300 2.996 x 104 ~ 3*100 = 300 Now you can order them any way you wish. Notice that here the numbers are wide apart, so approximation works well.